### v1

@@ Appendix 0.A Shading Module
Since we greatly improved the shading module, our implementation is quite different from the vanillar 3DGS, especially the backward part. In this section, we give the forward and backward propagation step by step so that everyone can learn the details and reproduce our shading module easily.
@@ 0.A.1 Forward
$\lambda_{1}=\frac{\text{Tr}(\hat{\bm{\Sigma}})+\sqrt{\text{Tr}(\hat{\bm{\Sigma}})^{2}-4\det(\hat{\bm{\Sigma}})}}{2},\quad$ $\lambda_{2}=\frac{\text{Tr}(\hat{\bm{\Sigma}})-\sqrt{\text{Tr}(\hat{\bm{\Sigma}})^{2}-4\det(\hat{\bm{\Sigma}})}}{2},$ (19)
$\hat{\bm{v}}_{1}=\begin{bmatrix}\hat{\Sigma}_{\text{12}}\\ \lambda_{1}-\hat{\Sigma}_{\text{11}}\end{bmatrix},\quad$ $\hat{\bm{v}}_{2}=\begin{bmatrix}\lambda_{2}-\hat{\Sigma}_{\text{22}}\\ \hat{\Sigma}_{\text{12}}\end{bmatrix},$
$\bm{v}_{1}=\begin{bmatrix}v_{1x}\\ v_{1y}\end{bmatrix}=\frac{\hat{\bm{v}}_{1}}{\|\hat{\bm{v}}_{1}\|}=\begin{bmatrix}\frac{\hat{\Sigma}_{\text{12}}}{\|\hat{\bm{v}}_{1}\|}\\ \frac{\lambda_{1}-\hat{\Sigma}_{\text{11}}}{\|\hat{\bm{v}}_{1}\|}\end{bmatrix},\quad$ $\bm{v}_{2}=\begin{bmatrix}v_{2x}\\ v_{2y}\end{bmatrix}=\frac{\hat{\bm{v}}_{2}}{\|\hat{\bm{v}}_{2}\|}=\begin{bmatrix}\frac{\lambda_{2}-\hat{\Sigma}_{\text{22}}}{\|\hat{\bm{v}}_{1}\|}\\ \frac{\hat{\Sigma}_{\text{12}}}{\|\hat{\bm{v}}_{1}\|}\end{bmatrix},$
@@ 0.A.2 Backward
$\frac{\partial S_{\sigma}(x)}{\partial x}$ $=\frac{\partial S\left(\frac{x}{\sigma}\right)}{\partial\frac{x}{\sigma}}\frac{\partial\frac{x}{\sigma}}{\partial x}=\left(1.6+0.021\cdot\left(\frac{x}{\sigma}\right)^{2}\right)S_{\sigma}\left(\frac{x}{\sigma}\right)\left[1-S_{\sigma}\left(\frac{x}{\sigma}\right)\right]\cdot\frac{1}{\sigma}$ (28)
$\frac{\partial S_{\sigma}(x)}{\partial\sigma}$ $=\frac{\partial S\left(\frac{x}{\sigma}\right)}{\partial\frac{x}{\sigma}}\frac{\partial\frac{x}{\sigma}}{\partial\sigma}=\left(1.6+0.021\cdot\left(\frac{x}{\sigma}\right)^{2}\right)S_{\sigma}\left(\frac{x}{\sigma}\right)\left[1-S_{\sigma}\left(\frac{x}{\sigma}\right)\right]\cdot-\frac{x}{\sigma^{2}}$
@@
It is quite difficult to directly express the above gradient. Still, we can use the chain rule to boil down the above results layer by layer to be available. We note that our key insight is to construct a new coordinate system using the mean vector $\hat{\bm{\mu}}$, eigenvalues $\{\lambda_{1},\lambda_{2}\}$ and eigenvectors $\{\bm{v}_{1},\bm{v}_{2}\}$. Therefore, we can use them as an intermediary to solve the final gradient by the chain rule. For the gradients of $\mathcal{I}_{g}^{\text{2D}}$ with respect to the mean vector $\hat{\bm{\mu}}$, according to Eq. 20 and Eq. 26, we have:
@@ Appendix 0.B Additional Results
In this section, we report more quantitative and qualitative results in detail.
@@ 0.B.3 Multi-scale Training and Multi-scale Testing on the Mip-NeRF 360 Dataset
PSNRnormal-$\uparrow$ bicycle flowers garden stump treehill
@@
SSIMnormal-$\uparrow$ bicycle flowers garden stump treehill
@@
LPIPSnormal-$\downarrow$ bicycle flowers garden stump treehill
@@
PSNRnormal-$\uparrow$ room counter kitchen bonsai
@@
SSIMnormal-$\uparrow$ room counter kitchen bonsai
@@
LPIPSnormal-$\downarrow$ room counter kitchen bonsai

### v2

@@ Appendix 0.A Shading Module
Since we greatly improved the shading module, our implementation is quite different from the vanilla 3DGS, especially the backward part. In this section, we give the forward and backward propagation step by step so that everyone can learn the details and reproduce our shading module easily.
@@ 0.A.1 Forward
$\lambda_{1}=\frac{\text{Tr}(\hat{\bm{\Sigma}})+\sqrt{\text{Tr}(\hat{\bm{\Sigma}})^{2}-4\det(\hat{\bm{\Sigma}})}}{2},$ $\lambda_{2}=\frac{\text{Tr}(\hat{\bm{\Sigma}})-\sqrt{\text{Tr}(\hat{\bm{\Sigma}})^{2}-4\det(\hat{\bm{\Sigma}})}}{2},$ (19)
$\hat{\bm{v}}_{1}=\begin{bmatrix}\hat{\Sigma}_{\text{12}}\\ \lambda_{1}-\hat{\Sigma}_{\text{11}}\end{bmatrix},$ $\hat{\bm{v}}_{2}=\begin{bmatrix}\lambda_{2}-\hat{\Sigma}_{\text{22}}\\ \hat{\Sigma}_{\text{12}}\end{bmatrix},$
$\bm{v}_{1}=\begin{bmatrix}v_{1x}\\ v_{1y}\end{bmatrix}=\frac{\hat{\bm{v}}_{1}}{\|\hat{\bm{v}}_{1}\|}=\begin{bmatrix}\frac{\hat{\Sigma}_{\text{12}}}{\|\hat{\bm{v}}_{1}\|}\\ \frac{\lambda_{1}-\hat{\Sigma}_{\text{11}}}{\|\hat{\bm{v}}_{1}\|}\end{bmatrix},$ $\bm{v}_{2}=\begin{bmatrix}v_{2x}\\ v_{2y}\end{bmatrix}=\frac{\hat{\bm{v}}_{2}}{\|\hat{\bm{v}}_{2}\|}=\begin{bmatrix}\frac{\lambda_{2}-\hat{\Sigma}_{\text{22}}}{\|\hat{\bm{v}}_{1}\|}\\ \frac{\hat{\Sigma}_{\text{12}}}{\|\hat{\bm{v}}_{1}\|}\end{bmatrix},$
@@ 0.A.2 Backward
$\frac{\partial S_{\sigma}(x)}{\partial x}$ $=\frac{\partial S\left(\frac{x}{\sigma}\right)}{\partial\frac{x}{\sigma}}\frac{\partial\frac{x}{\sigma}}{\partial x}=\left(1.6+0.21\cdot\left(\frac{x}{\sigma}\right)^{2}\right)S_{\sigma}\left(\frac{x}{\sigma}\right)\left[1-S_{\sigma}\left(\frac{x}{\sigma}\right)\right]\cdot\frac{1}{\sigma}$ (28)
$\frac{\partial S_{\sigma}(x)}{\partial\sigma}$ $=\frac{\partial S\left(\frac{x}{\sigma}\right)}{\partial\frac{x}{\sigma}}\frac{\partial\frac{x}{\sigma}}{\partial\sigma}=\left(1.6+0.21\cdot\left(\frac{x}{\sigma}\right)^{2}\right)S_{\sigma}\left(\frac{x}{\sigma}\right)\left[1-S_{\sigma}\left(\frac{x}{\sigma}\right)\right]\cdot-\frac{x}{\sigma^{2}}$
@@
It is quite difficult to express the above gradient directly. Still, we can use the chain rule to boil down the above results layer by layer to be available. We note that our key insight is to construct a new coordinate system using the mean vector $\hat{\bm{\mu}}$, eigenvalues $\{\lambda_{1},\lambda_{2}\}$ and eigenvectors $\{\bm{v}_{1},\bm{v}_{2}\}$. Therefore, we can use them as an intermediary to solve the final gradient by the chain rule. For the gradients of $\mathcal{I}_{g}^{\text{2D}}$ with respect to the mean vector $\hat{\bm{\mu}}$, according to Eq. 20 and Eq. 26, we have:
@@ Appendix 0.B Additional Results
In this section, we report more quantitative and qualitative results in detail. In addition, we provide the results of our method combined with 3D smoothing filter [36] for better study. According to the quantitative results of our experiments, the conclusion is that the 3D smoothing filter slightly improves the single-scale training and testing results. Still, it has no significant impact on the multi-scale training and testing results.
@@ 0.B.3 Multi-scale Training and Multi-scale Testing on the Mip-NeRF 360 Dataset
[…]
@@
Ours + 3D filter 25.32 27.50 29.04 29.88 21.79 24.04 26.49 28.52 27.19 29.52 31.14 32.08 26.80 28.15 29.87 31.48 22.72 24.33 25.97 27.49
[…]
@@
Ours + 3D filter 0.754 0.858 0.915 0.941 0.602 0.733 0.835 0.897 0.848 0.922 0.951 0.965 0.773 0.842 0.898 0.933 0.628 0.739 0.834 0.898
[…]
@@
[…]
@@
Ours + 3D filter 31.32 32.28 32.98 33.13 29.03 29.64 30.38 30.88 31.07 32.15 33.09 33.25 31.97 32.73 33.42 33.26
[…]
@@
Ours + 3D filter 0.915 0.946 0.964 0.971 0.903 0.924 0.943 0.955 0.924 0.951 0.967 0.973 0.939 0.958 0.969 0.972
[…]